A rectangular coil of n turns and area of cross section A , is held in a time varying magnetic field given by B=Bo sin wt , with the plane of the coil normal to the magnetic field . Deduce an expression for the emf induced in the coil . 

Grade:12

1 Answers

Snehal Jadhav
5 years ago
In this given question magnetic field(B) is variable.
Whereas, area(A) & angle (\theta) are constant.
 
\Phi = BAcos(\theta)
 
\frac{d\Phi }{dt}=\frac{dB}{dt}*Acos(\theta)
 
    = (w) Bo (-)sin(wt) A cos(\theta)
 
Emf = – \frac{d\Phi }{dt}
 
Emf = (w) Bo sin(wt) A cos(\theta)
       = Bo A (w) cos(\theta) sin(wt)